- Thanks to whoever helps.
Heather sold 3 packages of tulip bulbs and 2 packages of crocus bulbs for a total of $85. Jimmy sold 10 packages of tulip bulbs and 2 packages of crocus bulbs for a total of $190.

Answer by mananth(16946)

You must also ask a question.
Anyway
x cost of package of tulips
y cost of package of crocus
Total attended
3 x + 2 y = 85 .............1

10 x + 2 y = 190 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-3 x -2 y = -85
10 x + 2 y = 190
Add the two equations
7 x = 105
/ 7
x = 15
plug value of x in (1)
3 x + 2 y = 85
45 + 2 y = 85
2 y = 85 -45.00
2 y = 40
y = 20
$15 cost of package of tulips
$20 cost of package of crocus
m.ananth@hotmail.ca